y=3x
y-4x=10
solve for x

Answer by mananth(12270)

y=3x
y-4x=10
subtract the second equation from the first.
y-(y-4x) =3x-10
y-y+4x=3x-10
4x=3x-10
4x-3x=10
x=10
.
y=30